# Learning when to rank: Estimation of partial rankings from sparse, noisy comparisons

### v2

@@ Abstract
A common task arising in various domains is that of ranking items based on the outcomes of pairwise comparisons, from ranking players and teams in sports to ranking products or brands in marketing studies and recommendation systems. Statistical inference-based methods such as the Bradley-Terry model, which extract rankings based on an underlying generative model of the comparison outcomes, have emerged as flexible and powerful tools to tackle the task of ranking in empirical data. In situations with limited and/or noisy comparisons, it is often challenging to confidently distinguish the performance of different items based on the evidence available in the data. However, existing inference-based ranking methods overwhelmingly choose to assign each item to a unique rank or score, suggesting a meaningful distinction when there is none. Here, we address this problem by developing a principled Bayesian methodology for learning partial rankings—rankings with ties—that distinguishes among the ranks of different items only when there is sufficient evidence available in the data. Our framework is adaptable to any statistical ranking method in which the outcomes of pairwise observations depend on the ranks or scores of the items being compared. We develop a fast agglomerative algorithm to perform Maximum A Posteriori (MAP) inference of partial rankings under our framework and examine the performance of our method on a variety of real and synthetic network datasets, finding that it frequently gives a more parsimonious summary of the data than traditional ranking, particularly when observations are sparse.
@@ I Introduction
Given the prominence of rankings across domains, it has long been of interest to develop algorithms that automatically rank items based on some set of data about each item bradley1952rank ; mallows1957non ; luce1959individual ; plackett1975analysis ; langville2012s . An extensively studied subset of ranking problems is that of ranking from pairwise comparisons whelan2017prior ; davidson1973bayesian ; langville2012s ; david1963method ; bradley1952rank ; jerdee2024luck , where rankings are inferred from comparisons among pairs of entities. In this approach, the outcome of a single interaction between two items—for example, a win or loss in sports or one product being preferred over another—serves as the basis for the ranking. This approach allows one to infer rankings in situations where underlying ratings of the entities are unavailable, unobserved, or difficult to compute. For example, instead of trying to assess the absolute skill of individual tennis players, we can infer their relative rankings from their head-to-head match results, assuming that matches are more often won by the better player.
@@
Since pairwise comparisons inherently involve some degree of uncertainty due to randomness, noise, or incomplete information in the observed outcomes, pairwise ranking models are generally formulated as probabilistic models whose parameters can then be estimated, for example, via Maximum Likelihood Estimation (MLE) or Maximum A Posteriori (MAP) inference bradley1952rank ; elo1978rating ; herbrich2006trueskill ; de2018physical ; peixoto2022ordered . It is also useful to model pairwise comparison data as a network in which items are nodes and each comparison is an edge between its participating nodes, pointing from the winner of the corresponding match/preferred item to the loser of the match/non-preferred item. Multiple edges may exist between a pair of nodes if multiple comparisons were performed, each edge providing evidence for the relative ranking of the two items on its endpoints.
@@
In the classic Bradley-Terry (BT) model bradley1952rank , the outcome of a comparison between two nodes is modeled as a Bernoulli random variable with a probability that depends on the latent scores associated with the items. After inference with the model, one obtains a set of continuous scores that can be ordered to obtain a final ranking. The BT model has inspired numerous extensions agresti2012categorical ; jerdee2024luck to more flexibly model variations in match outcomes, including the possibility of ties in the match outcomes rao1967ties ; davidson1970extending ; yan2016ranking ; baker2021modifying (rather than ties in the final rankings as we study here). Another popular inference-based method for ranking from pairwise comparisons is SpringRank de2018physical , which ranks nodes by finding the ground state of a physical system which models the comparison network as a set of directed springs and has also inspired numerous generalizations iacovissi2022interplay ; della2024model . These methods can parsimoniously model heterogeneity and noise in match outcomes but do not allow for the possibility of ties in the inferred rankings and, as such, are unable to determine when there is enough structure in the comparison data to justify distinguishing the rankings of different items. (In principle, two players can have identical scores under such methods, but this is rarely ever achieved in practice numerically.) Such methods can therefore infer rankings that overfit the data when limited observations are available, resulting in rankings that are highly sensitive to small changes in outcomes. Several methods have been proposed to reduce this overfitting and promote ties in the inferred rankings (see Sec. II.2). However, these models require extensive parameter tuning, a priori knowledge about the number and size of the groups being inferred, or incorporate information other than the outcomes of the pairwise comparisons, making them difficult to adapt to a broad range of existing ranking methods.
@@
The paper is organized as follows. In Sec. II.1 and Sec. II.2, we introduce the BT model and review literature relating to partial ranking inference. In Sec. II.3 and Sec. II.4 and we introduce our partial rankings model and the agglomerative algorithm we use to perform MAP inference with our model. In Sec. III.1 and Sec. III.2, we apply our method to a wide range of synthetically generated datasets as well as a corpus of empirical networks, finding that partial rankings can provide a more parsimonious description of the data in cases where the networks of outcomes are not sufficiently dense. In Sec. III.3, we focus on a network of faculty hiring among computer science departments in the U.S. and show that our algorithm can be used to extract meaningful ambiguities in the rankings of these departments. We finalize in Sec. IV with our conclusions.
@@ II.1 Bayesian Ranking and the Bradley-Terry Model
For illustrative purposes, it will be convenient to discuss ranking in the context of matches among individuals—for example, tennis matches among tennis players. But all upcoming discussion can be equivalently framed in the context of comparing two items, in which case each match is a comparison (e.g. from a respondent in a marketing survey or a consumer purchasing one of multiple products), and the winner of the match is the preferred item.
Let $\bm{W}$ be an $N\times N$ matrix of match outcomes for a set of $N$ competing individuals (nodes) such that $w_{ij}$ is the number of matches in which node $i$ beat node $j$. For simplicity, we will assume that ties in the matches are not allowed. The problem we want to solve is that of inferring a ranking $\bm{r}=[r_{1},...,r_{N}]$ of the $N$ individuals such that $r_{i}<r_{j}$ indicates a higher probability of $i$ beating $j$ than $j$ beating $i$. Given the inherently probabilistic nature of match outcomes, a common approach is to assign a latent real-valued score $s_{i}\in\mathbb{R}$ to each player such that the probability $p_{ij}$ that player $i$ beats player $j$ is given by some function of the difference of their scores: $p_{ij}=f(s_{i}-s_{j})$. For $f(s)$ to be a suitable scoring function, it must satisfy a set of constraints jerdee2024luck : (1) It must be a monotonically increasing function of $s$, ensuring that stronger players are assigned a higher probability of winning; (2) It must be bounded within $[0,1]$ as it must represent a valid probability; (3) It must be antisymmetric around zero, satisfying $f(-s)=1-f(s)$ to ensure that the probability of losing is one minus the probability of winning. While these requirements do not uniquely determine $f(s)$, a commonly adopted choice is the logistic function, $f(s)=1/(1+e^{-s})$, which gives
@@
For convenience, one generally introduces the quantities $\pi_{i}=e^{s_{i}}$, which, keeping with previous literature, we shall call the player strengths zermelo1929berechnung ; newman2023efficient . Eq. 1 can then be written as
@@
Differentiating the log-likelihood (i.e. the logarithm of Eq. 3) with respect to the $\pi_{i}$ and equating the result to zero, we obtain the following set of self-consistent equations:
@@
which have all been proven to converge to the same MLE solution (if one exists) for $0\leq\alpha\leq 1$. Furthermore, it has been shown that the convergence of this family of algorithms becomes monotonically slower with increasing $\alpha$, with $\alpha=0$ converging fastest and Zermelo’s algorithm (corresponding to the $\alpha=1$ case) being the slowest to converge. Eq. 5 was first derived in newman2023efficient , and we use its $\alpha=0$ realization throughout this work.
The MLE approach described above presents a series of drawbacks. First, we note that Eq. 2, and therefore Eq. 3, is invariant under a constant rescaling of all player strengths, meaning that player scores are not uniquely identifiable. To resolve this ambiguity, one generally imposes a suitable normalization. A common choice is to fix the average player score to zero
@@
More problematic is the fact that Eq. 5—and all MLE methods for that matter—have been shown to converge only under the stringent condition that the underlying network of matches is strongly connected, i.e. there exists a directed path between any pair of nodes in the network zermelo1929berechnung ; ford1957solution ; hunter2004mm . If this is not the case, then no maximum of the likelihood function exists, and the scores in Eq. 5 diverge.
@@
Eq. 9 removes the invariance under a rescaling of the scores. Furthermore, it also prevents the scores from diverging. As pointed out in whelan2017prior , the prior for an individual $i$ can be written as
@@
which corresponds to the probability that player $i$ plays two matches against the average player, winning one and losing the other. Eq. 9 can then be seen as the likelihood of a BT model where two pseudo-games have been added for each player, one won and one lost. The resulting network of interactions is now strongly connected, ensuring that the maximum of the posterior always exists, irrespective of whether the original interaction network was fully connected or not.
@@
which is the generalization of Eq. 5 (with $\alpha=0$) to the MAP setting and can be solved in the same way as for the MLEs. The primary advantage of Eq. 12 over the MLE updates is that the iterative updates described by Eq. 12 are guaranteed to converge, and do so quite efficiently in practice newman2023efficient . The inferred strengths are then converted into an implied ranking $\bm{r}$ by ordering the nodes with respect to the inferred strengths $\bm{\hat{\pi}}$.
@@ II.2 Partial Rankings
Many of these issues can be resolved by resorting to Bayesian approaches. However, these approaches generally require Monte Carlo estimation, which can be slow for estimation davidson1973bayesian ; caron2012efficient . In pearce2024bayesian , Pearce and Erosheva propose a Bayesian method based on the use of a spike-and-slab prior mitchell1988bayesian ; george1997approaches ; ishwaran2005spike to induce parameter fusion. Unlike other methodologies, Pearce and Erosheva’s model requires neither the specification of the number or sizes of the rank cluster nor the tuning of lasso-type penalties, while exploiting conjugacy relationships for estimation of the full posterior. However, this model still requires the selection and tuning of a series of hyperparameters, which influence how aggressively the model fuses the rankings. Furthermore, the particular choice of priors, while facilitating computation and providing strong regularization, might imbue unwarranted structure into the model, particularly in the sparse case, where the data might be insufficient to allow the prior to wash away.
@@ II.3 Bayesian Nonparametric Model
Given Eq. II.3, we now need to define the functional forms of the likelihood and priors. The easiest choice is to use the BT likelihood of Eq. 3, although any choice of likelihood is possible to incorporate into this MAP estimation objective so long as it is a function of latent strengths—i.e., $P(\bm{W}|\bm{\pi})$—or ranks—i.e., $P(\bm{W}|\bm{r})$ (the latter allowing us to remove the prior $P(\bm{\pi}|\bm{r})$). This flexibility allows for the inclusion of a wide range of ranking objectives including extensions of BT models jerdee2024luck and SpringRank de2018physical . For our experiments in Sec. III, we will proceed with using the BT likelihood of Eq. 3 for subsequent analysis, which facilitates an efficient optimization scheme and a simple comparison with the original BT model.
@@
    Draw the histogram $\bm{n}=[n_{1},...,n_{R}]$ of the number of nodes of each rank uniformly from the set of all histograms compatible with the number of ranks $R$. This amounts to selecting a list of $R$ positive integer values that sum to $N$. As there are ($N-1\choose R-1$ such lists, each draw occurs with probability
@@
    Draw the ranks $\bm{r}$ uniformly over the set of rankings compatible with the rank counts $\bm{n}$. Since there are ($N\choose n_{1},...,n_{R}$ possible partitions of $N$ objects into $R$ groups such that each group $r$ has size $n_{r}$, the total probability of this step is given by
@@
A reasonable choice for the prior $P(\bm{\pi}|\bm{r})$ is a hierarchical prior consisting of two steps: (1) Draw the set of unique strengths $\bm{\sigma}=\{\sigma_{1},\cdots,\sigma_{R}\}$ based on the number of unique ranks $R$; (2) Draw the individual strengths $\bm{\pi}$ using the deterministic prior $P(\bm{\pi}|\bm{\sigma})=\prod_{i=1}^{N}\delta(\pi_{i},\sigma_{r_{i}})$. Combining these gives the prior
@@
denoting the number of edges going from nodes of rank $r$ to nodes of rank $r^{\prime}$ (i.e. the number of times nodes of rank $r$ beat nodes of rank $r^{\prime}$). By minimizing Eq. II.3, we can determine the optimal ranks $\bm{\hat{r}}$ and the corresponding optimal strengths $\bm{\hat{\sigma}}=\{\hat{\sigma}_{\hat{r}_{i}}\}_{i=1}^{N}$ that best fit the observed data. It is important to emphasize that, although this work focuses primarily on MAP estimates of the strengths and rankings, Eq. II.3 defines a well-posed (negative) log-posterior distribution. This posterior can be explored using sampling techniques such as Markov Chain Monte Carlo (MCMC), enabling estimation of the full posterior distribution and associated uncertainty quantification rather than relying solely on point estimates.
We can see that the first term in Eq. II.3 is a constant independent of the underlying node ranking. The second and third terms synergistically penalize having a large number of rankings as they will increase $\mathcal{L}(\bm{r},\bm{\sigma})$. Indeed, it is easy to prove that the sum of these two terms
@@
By taking the logarithm of Eq. 26, we can rewrite the posterior odds ratio as

### v3

@@ Abstract
Ranking items based on pairwise comparisons is common, from using match outcomes to rank sports teams to using purchase or survey data to rank consumer products. Statistical inference-based methods such as the Bradley-Terry model, which extract rankings based on an underlying generative model, have emerged as flexible and powerful tools to tackle ranking in empirical data. In situations with limited and/or noisy comparisons, it is often challenging to confidently distinguish performance of different items based on evidence available in the data. However, most inference-based ranking methods choose to assign each item to a unique rank or score, suggesting a meaningful distinction when there is none. Here, we develop a principled nonparametric Bayesian method, adaptable to any statistical ranking method, for learning partial rankings (rankings with ties) that distinguishes among the ranks of different items only when there is sufficient evidence available in the data. We develop a fast agglomerative algorithm to perform Maximum A Posteriori (MAP) inference of partial rankings under our framework and examine the performance of our method on a variety of real and synthetic network datasets, finding that it frequently gives a more parsimonious summary of the data than traditional ranking, particularly when observations are sparse.
@@ I Introduction
Given the prominence of rankings across domains, it has long been of interest to develop algorithms that automatically rank items based on some set of data about the items and their associations bradley1952rank ; mallows1957non ; luce1959individual ; plackett1975analysis ; langville2012s . An extensively studied subset of ranking problems is that of ranking from pairwise comparisons whelan2017prior ; davidson1973bayesian ; langville2012s ; david1963method ; bradley1952rank ; jerdee2024luck , where rankings are inferred from comparisons among pairs of entities. In this approach, the outcome of a single interaction between two items—for example, a win or loss in sports or one product being preferred over another—serves as the basis for the ranking. This approach allows one to infer rankings in situations where underlying ratings of the entities are unavailable, unobserved, or difficult to compute. For example, instead of trying to assess the absolute skill of individual tennis players, we can infer their relative rankings from their head-to-head match results, assuming that matches are more often won by the better player.
@@
Since pairwise comparisons inherently involve some degree of uncertainty due to randomness, noise, or incomplete information in the observed outcomes, pairwise ranking models are generally formulated as probabilistic models whose parameters can then be estimated, for example, via Maximum Likelihood Estimation (MLE) or Maximum A Posteriori (MAP) inference bradley1952rank ; elo1978rating ; herbrich2006trueskill ; de2018physical ; peixoto2022ordered . It is also useful to model pairwise comparison data as a network in which items are nodes and each comparison is an edge between its participating nodes, pointing from the winner of the corresponding match/preferred item to the loser of the match/non-preferred item. Multiple edges may exist between a pair of nodes if multiple comparisons were performed, each edge providing evidence for the relative ranking of the two items on its endpoints. Therefore, we will use the network-centric terminology “nodes” interchangeably with the entities being ranked (e.g. players, teams, etc) and “edges” interchangeably with the pairwise comparisons used for ranking (e.g. matches, hiring flows, etc).
@@
In the classic Bradley-Terry (BT) model bradley1952rank , the outcome of a comparison between two nodes is modeled as a Bernoulli random variable with a probability that depends on the latent scores associated with the items. After inference with the model, one obtains a set of continuous scores that can be ordered to obtain a final ranking. The BT model has inspired numerous extensions agresti2012categorical ; jerdee2024luck to more flexibly model variations in match outcomes, including the possibility of ties in the match outcomes rao1967ties ; davidson1970extending ; yan2016ranking ; baker2021modifying (rather than ties in the final rankings as we study here). Another popular inference-based method for ranking from pairwise comparisons is SpringRank de2018physical , which ranks nodes by finding the ground state of a physical system which models the comparison network as a set of directed springs and has also inspired numerous generalizations iacovissi2022interplay ; della2024model . These methods can parsimoniously model heterogeneity and noise in match outcomes but do not allow for the possibility of ties in the inferred rankings and, as such, are unable to determine when there is enough structure in the comparison data to justify distinguishing the rankings of different items. (In principle, two nodes can have identical scores under such methods, but this is rarely ever achieved in practice numerically.) Such methods can therefore infer rankings that overfit the data when limited observations are available, resulting in rankings that are highly sensitive to small changes in outcomes. Several methods have been proposed to reduce this overfitting and promote ties in the inferred rankings (see Sec. II.2). However, these models often require extensive parameter tuning, a priori knowledge about the number and size of the groups being inferred, or incorporate information other than the outcomes of the pairwise comparisons, making them difficult to adapt to a broad range of existing ranking methods.
@@
The paper is organized as follows. In Sec. II.1 and Sec. II.2, we introduce the BT model and review literature relating to partial ranking inference. In Sec. II.3 and Sec. II.4 and we introduce our partial rankings model and the agglomerative algorithm we use to perform MAP inference with our model. In Sec. III.1 and Sec. III.2, we apply our method to a wide range of synthetically generated datasets as well as a corpus of empirical networks, finding that partial rankings can provide a more parsimonious description of the data in cases where the networks of outcomes are not sufficiently dense. In Sec. III.3, we focus on a network of faculty hiring among computer science departments in the U.S. and show that our algorithm can be used to extract ambiguities in the rankings of these departments. We finalize in Sec. IV with our conclusions.
@@ II.1 Bayesian Ranking and the Bradley-Terry Model
For illustrative purposes, it will be convenient to discuss ranking in the context of matches among individuals—for example, tennis matches among tennis players. But all upcoming discussion can be equivalently framed in the context of comparing two items, in which case each match is a comparison (e.g. from a respondent in a marketing survey or a consumer purchasing one of multiple products), and the winner of the match is the preferred item. All such entities can be captured by the nodes in a network of the pairwise comparisons.
Let $\bm{W}$ be an $N\times N$ matrix of match outcomes for a set of $N$ competing individuals (nodes) such that $w_{ij}$ is the number of matches in which node $i$ beat node $j$. For simplicity, we will assume that ties in the matches are not allowed. The problem we want to solve is that of inferring a ranking $\bm{r}=[r_{1},...,r_{N}]$ of the $N$ individuals such that $r_{i}<r_{j}$ indicates a higher probability of $i$ beating $j$ than $j$ beating $i$. Given the inherently probabilistic nature of match outcomes, a common approach is to assign a latent real-valued score $s_{i}\in\mathbb{R}$ to each player such that the probability $p_{ij}$ that player $i$ beats player $j$ is given by some function of the difference of their scores: $p_{ij}=f(s_{i}-s_{j})$. For $f(s)$ to be a suitable scoring function, it must satisfy a set of constraints jerdee2024luck : (1) It must be a monotonically increasing function of $s$, ensuring that stronger players are assigned a higher probability of winning; (2) It must be bounded within $[0,1]$ as it must represent a valid probability; (3) It must be antisymmetric around zero, satisfying f(s)=1f(s)f(-s)=1-f(s) to ensure that the probability of losing is one minus the probability of winning. While these requirements do not uniquely determine $f(s)$, a commonly adopted choice is the logistic function, $f(s)=1/(1+e^{-s})$, which gives
@@
For convenience, one generally introduces the (non-negative) quantities $\pi_{i}=e^{s_{i}}$, which, keeping with previous literature, we shall call the player strengths zermelo1929berechnung ; newman2023efficient . Eq. (1) can then be written as
@@
Differentiating the log-likelihood (i.e. the logarithm of Eq. (3)) with respect to the $\pi_{i}$ and equating the result to zero, we obtain the following set of self-consistent equations:
@@
which have all been proven to converge to the same MLE solution (if one exists) for $0\leq\alpha\leq 1$. Furthermore, it has been shown that the convergence of this family of algorithms becomes monotonically slower with increasing $\alpha$, with $\alpha=0$ converging fastest and Zermelo’s algorithm (corresponding to the $\alpha=1$ case) being the slowest to converge. Eq. (5) was first derived in newman2023efficient , and we use its $\alpha=0$ realization throughout this work.
The MLE approach described above presents a series of drawbacks. First, we note that Eq. (2), and therefore Eq. (3), is invariant under a constant rescaling of all player strengths, meaning that player scores are not uniquely identifiable. To resolve this ambiguity, one generally imposes a suitable normalization. A common choice is to fix the average player score to zero
@@
More problematic is the fact that Eq. (5)—and all MLE methods for that matter—have been shown to converge only under the stringent condition that the underlying network of matches is strongly connected, i.e. there exists a directed path between any pair of nodes in the network zermelo1929berechnung ; ford1957solution ; hunter2004mm . If this is not the case, then no maximum of the likelihood function exists, and the scores in Eq. (5) diverge.
@@
Eq. (9) removes the invariance under a rescaling of the scores. Furthermore, it also prevents the scores from diverging. As pointed out in whelan2017prior , the prior for an individual $i$ can be written as
@@
which corresponds to the probability that player $i$ plays two matches against the average player, winning one and losing the other. Eq. (9) can then be seen as the likelihood of a BT model where two pseudo-games have been added for each player, one won and one lost. The resulting network of interactions is now strongly connected, ensuring that the maximum of the posterior always exists, irrespective of whether the original interaction network was fully connected or not.
@@
which is the generalization of Eq. (5) (with $\alpha=0$) to the MAP setting and can be solved in the same way as for the MLEs. The primary advantage of Eq. (12) over the MLE updates is that the iterative updates described by Eq. (12) are guaranteed to converge, and do so quite efficiently in practice newman2023efficient . The inferred strengths are then converted into an implied ranking $\bm{r}$ by ordering the nodes with respect to the inferred strengths $\bm{\hat{\pi}}$.
@@ II.2 Partial Rankings
Many of these issues can be resolved by resorting to Bayesian approaches. However, these approaches generally require Monte Carlo estimation, which can be slow for estimation davidson1973bayesian ; caron2012efficient . In pearce2024bayesian , Pearce and Erosheva propose the Rank-Clustered Bradley-Terry-Luce (RC-BTL) model, a Bayesian method based on the use of a spike-and-slab prior mitchell1988bayesian ; george1997approaches ; ishwaran2005spike to induce parameter fusion. Unlike other methodologies, Pearce and Erosheva’s model requires neither the specification of the number or sizes of the rank cluster nor the tuning of lasso-type penalties, while exploiting conjugacy relationships for estimation of the full posterior. However, this model still requires the selection and tuning of a series of hyperparameters, among which is setting the expected number of rank clusters. Although this choice can be advantageous, as it allows the user to tune the desired level of rank clustering, poor choices of this parameter could potentially introduce unwanted biases in the inferred results.
@@ II.3 Bayesian Nonparametric Model
Given Eq. (II.3), we now need to define the functional forms of the likelihood and priors. The easiest choice is to use the BT likelihood of Eq. (3), although any choice of likelihood is possible to incorporate into this MAP estimation objective so long as it is a function of latent strengths—i.e., $P(\bm{W}|\bm{\pi})$—or ranks—i.e., $P(\bm{W}|\bm{r})$ (the latter allowing us to remove the prior $P(\bm{\pi}|\bm{r})$). This flexibility allows for the inclusion of a wide range of ranking objectives, including extensions of BT models jerdee2024luck and SpringRank de2018physical . For our experiments in Sec. III, we will proceed with using the BT likelihood of Eq. (3) for subsequent analysis, which facilitates an efficient optimization scheme and a simple comparison with the original BT model.
@@
    Draw the histogram $\bm{n}=[n_{1},...,n_{R}]$ of the number of nodes of each rank uniformly from the set of all histograms compatible with the number of ranks $R$. This amounts to selecting a list of $R$ positive integer values that sum to $N$. As there are $N-1\choose R-1$ such lists, each draw occurs with probability
@@
    Draw the ranks $\bm{r}$ uniformly over the set of rankings compatible with the rank counts $\bm{n}$. Since there are $N\choose n_{1},...,n_{R}$ possible partitions of $N$ objects into $R$ groups such that each group $r$ has size $n_{r}$, the total probability of this step is given by
@@
A reasonable choice for the prior $P(\bm{\pi}|\bm{r})$ is a hierarchical prior consisting of two steps: (1) Draw the set of (non-negative) unique strengths $\bm{\sigma}=\{\sigma_{1},\cdots,\sigma_{R}\}$ based on the number of unique ranks $R$; (2) Draw the individual strengths $\bm{\pi}$ using the deterministic prior $P(\bm{\pi}|\bm{\sigma})=\prod_{i=1}^{N}\delta(\pi_{i},\sigma_{r_{i}})$. Combining these gives the prior
@@
denoting the number of edges going from nodes of rank $r$ to nodes of rank $r^{\prime}$ (i.e. the number of times nodes of rank $r$ beat nodes of rank $r^{\prime}$).
The generative process described above, and the associated objective, are similar to those of the RC-BTL model introduced by Pearce and Erosheva pearce2024bayesian , which places a Gamma prior on the unique cluster strengths. The main difference between our model and RC-BTL lies in the prior over partitions. In our Partial Ranking (PR) formulation, players are assigned to groups via a hierarchy of uniform priors over the number, sizes, and compositions of the clusters. By contrast, RC-BTL employs a Poisson prior on the number of clusters and assigns equal probability to all partitions with the same $R$, regardless of cluster sizes. As a result, PR and RC-BTL encode different structural biases over the space of partitions. In RC-BTL, the indifference with respect to cluster sizes simplifies posterior sampling, since all partitions with the same number of groups are equally likely, but it also implicitly favors balanced groups: if partitions are drawn uniformly conditional on $R$, the overwhelming majority will allocate players approximately equally across groups peixoto2019bayesian ; peixoto2017nonparametric . Depending on the amount of available data and prior knowledge about the items being ranked, this may introduce artificial structure into the inferred rankings. In contrast, the PR formulation explicitly samples the group sizes as part of the prior hierarchy, remaining agnostic about balance until the data are observed. For a more detailed comparison between the PR and RC-BTL models see Appendix F.
By minimizing Eq. (II.3), we can determine the optimal ranks $\bm{\hat{r}}$ and the corresponding optimal strengths $\bm{\hat{\sigma}}=\{\hat{\sigma}_{\hat{r}_{i}}\}_{i=1}^{N}$ that best fit the observed data. It is important to emphasize that, although this work focuses primarily on MAP estimates of the strengths and rankings, Eq. (II.3) defines a well-posed (negative) log-posterior distribution. This posterior can be explored using sampling techniques such as Markov Chain Monte Carlo (MCMC), enabling estimation of the full posterior distribution and associated uncertainty quantification rather than relying solely on point estimates.
We can see that the first term in Eq. (II.3) is a constant independent of the underlying node ranking. The second and third terms synergistically penalize having a large number of rankings as they will increase $\mathcal{L}(\bm{r},\bm{\sigma})$. Indeed, it is easy to prove that the sum of these two terms
@@
By taking the logarithm of Eq. (26), we can rewrite the posterior odds ratio as